# The electronic and magnetic structures of bilayer La3Ni2O7 at ambient pressure

### v2

@@ Abstract
.1 Exchange-correlation functional, GGA, and Hybrid functional
DFT is a powerful computational method to investigate the electronic structure of solids. The basic computational methods behind DFT calculations are the Kohn-Sham equations, which provide a systematic way of mapping the many-body system onto a single-body problem martin2020electronic . More precisely, the total energy of the system in Kohn-Sham theory is written as a functional of electron density $\rho(\mathbf{r})$,
@@
The $T_{s}[\rho]$ is the Kohn-Sham kinetic energy, $v_{ext}(\mathbf{r})$ is the external potential like the electron-nuclei interaction, and $E_{H}[\rho]$ is the Hatree energy. The $E_{xc}[\rho]$ is the exchange-correlation functional. $T_{s}$, $v_{ext}$, $E_{H}[\rho]$ can be straightforwardly calculated. The only unknown in Kohn-Sham theory is the exchange-correlation functional $E_{xc}$ martin2020electronic ; Jacob_ladder . Developing a precise $E_{xc}$ is one central goal in the DFT framework. Hence, implementing a precise $E_{xc}$ is a crucial step for band structure calculation. The most commonly used $E_{xc}$ in solid-state physics is the GGA or its PBE version perdew1996generalized . The GGA band structures of LNO in the LP and HP phases are plotted in Fig. S1. In both cases, the bonding $d_{z}$ bands cross the Fermi level, which is due to the underestimated exchange correlations.
Going beyond GGA, a more advanced exchange-correlation functional $E_{xc}$ is the hybrid functional. The exchange part or correlation part of $E_{xc}$ is a linear combination of the Hatree-Fock (HF) exchange and semilocal potential such as GGA. The hybrid functionals can be divided into two families according to the interelectronic range at which the HF exchange is applied: screened (range-separated) and unscreened. The screened hybrid functional like sophisticated HSE06 which mixes GGA and HF in the short-range exchange energy is often used in solid state calculations. It provides great accuracy for band gaps exchange_correlation . There are also other unscreened hybrid functionals like the PBE0 perdew1996rationale ; ernzerhof1999assessment ; adamo1999toward and B3LYP stephens1994ab , which have slower convergence than HSE06 and also been tested. PBE0 is similar to HSE06 but mixes both the short-range and the long-range contributions. B3LYP mixes LDA, HF, and Becke88 in the exchange energy. The correlation energy is also a mixture of LYP and VWN3. The PBE0 and B3LYP band structures are shown in Fig. S2, which are similar to HSE06 in the main text.
@@
Although the hybrid functional provides an accurate description of band structure, the local Hubbard interactions beyond DFT should be treated using other many-body techniques. One commonly used method is DMFT. In the main text, we presented a DMFT study of the spectral function in the LP phase. The basic assumption of DMFT is that the self-energy only depends on the frequency $\Sigma(\omega)$. We use the open-source TRIQS DMFT package triqs and its CTQMC solver ctqmc . In the LP phase, there are four Ni atoms. To simplify our discussion, we treat them equally in the DMFT calculation. During each DMFT self-consistent calculation, a $40\times 40$ Brillouin zone mesh is used for the lattice Green’s function calculation.
@@
Therefore, an immediate question is whether using only the traditional GGA functional, after considering many-body interactions, can yield band structures consistent with experiments. Hence, we also performed DMFT calculations using a TB model fitted from previous band calculations yaodx with the GGA functional, as shown in Fig. S3(a). It is evident that even with interactions included, the $d_{z^{2}}$ bonding state is not pushed below the Fermi level.
@@
The DFT+U is another convenient approach martin2020electronic . However, DFT+U only shifts the crystal field splitting without renormalizing the hopping parameters or the bandwidths. In previous works, it was proposed that DFT+U (using GGA functional) can capture the ARPES measurements of the band dispersion in ambient pressure zhouxj . We also calculate the band structure of the LP phase using the simplified rotation invariant approach to the DFT+U introduced by Dudarev $et$ $al.$ dudarev1998electron based on GGA functional. As shown in Fig. S3(b), the DFT+U results with $U$=4eV show significant deviation from the ARPES measurements, particularly in terms of the renormalization effect, despite the band positions being nearly correct.
@@
Figure S3: (a) The spectral function $A(k,\omega)$ calculated using DMFT at $U$ = $5$ eV, $J_{H}$ = $0.6$ eV, and $\beta=50$ based on TB bands fitted from DFT bands of ordinary GGA functional. We want to add a note here. The $U/J_{H}$ here is little smaller than the main text owing to the smaller spectra weight at large U. The conclusion remains valid. (b) The DFT+U band dispersion using GGA functional at $U=4$ eV. The calculated band dispersion cannot describe the experiment data (red crosses) owing to the band position being incorrect in (a) and missing additional renormalization effects in (b).
.3 Hamiltonian and Tight-binding parameters

### v3

@@ VI SUMMARY
VI.1 Exchange-correlation functional, GGA, and Hybrid functional
DFT is a powerful computational method to investigate the electronic structure of solids. The basic computational methods behind DFT calculations are the Kohn-Sham equations, which provide a systematic way of mapping the many-body system onto a single-body problem Martin (2020). More precisely, the total energy of the system in Kohn-Sham theory is written as a functional of electron density $\rho(\mathbf{r})$,
@@ VI.1 Exchange-correlation functional, GGA, and Hybrid functional
The $T_{s}[\rho]$ is the Kohn-Sham kinetic energy, $v_{ext}(\mathbf{r})$ is the external potential like the electron-nuclei interaction, and $E_{H}[\rho]$ is the Hatree energy. The $E_{xc}[\rho]$ is the exchange-correlation functional. $T_{s}$, $v_{ext}$, $E_{H}[\rho]$ can be straightforwardly calculated. The only unknown in Kohn-Sham theory is the exchange-correlation functional $E_{xc}$ Martin (2020); Perdew and Schmidt (2001). Developing a precise $E_{xc}$ is one central goal in the DFT framework. Hence, implementing a precise $E_{xc}$ is a crucial step for band structure calculation. The most commonly used $E_{xc}$ in solid-state physics is the GGA or its PBE version Perdew et al. (1996a). The GGA band structures of LNO in the LP and HP phases are plotted in Fig. S1. In both cases, the bonding $d_{z}$ bands cross the Fermi level, which is due to the underestimated exchange correlations.
Going beyond GGA, a more advanced exchange-correlation functional $E_{xc}$ is the hybrid functional. The exchange part or correlation part of $E_{xc}$ is a linear combination of the Hatree-Fock (HF) exchange and semilocal potential such as GGA. The hybrid functionals can be divided into two families according to the interelectronic range at which the HF exchange is applied: screened (range-separated) and unscreened. The screened hybrid functional like sophisticated HSE06 which mixes GGA and HF in the short-range exchange energy is often used in solid state calculations. It provides great accuracy for band gaps Borlido et al. (2019). There are also other unscreened hybrid functionals like the PBE0 Perdew et al. (1996b); Ernzerhof and Scuseria (1999); Adamo and Barone (1999) and B3LYP Stephens et al. (1994), which have slower convergence than HSE06 and also been tested. PBE0 is similar to HSE06 but mixes both the short-range and the long-range contributions. B3LYP mixes LDA, HF, and Becke88 in the exchange energy. The correlation energy is also a mixture of LYP and VWN3. The PBE0 and B3LYP band structures are shown in Fig. S2, which are similar to HSE06 in the main text.
@@ VI.2 DMFT and DFT+U
Although the hybrid functional provides an accurate description of band structure, the local Hubbard interactions beyond DFT should be treated using other many-body techniques. One commonly used method is DMFT. In the main text, we presented a DMFT study of the spectral function in the LP phase. The basic assumption of DMFT is that the self-energy only depends on the frequency $\Sigma(\omega)$. We use the open-source TRIQS DMFT package Parcollet et al. (2015) and its CTQMC solver Seth et al. (2016). In the LP phase, there are four Ni atoms. To simplify our discussion, we treat them equally in the DMFT calculation. During each DMFT self-consistent calculation, a $40\times 40$ Brillouin zone mesh is used for the lattice Green’s function calculation.
@@
Therefore, an immediate question is whether using only the traditional GGA functional, after considering many-body interactions, can yield band structures consistent with experiments. Hence, we also performed DMFT calculations using a TB model fitted from previous band calculations Luo et al. (2023) with the GGA functional, as shown in Fig. S3(a). It is evident that even with interactions included, the $d_{z}$ bonding state is not fully pushed below the Fermi level. From Fig. S3(c), by examining the real part of the self-energy and using the method mentioned in the main text, we can also determine that the band top of the $d_{z}$ bonding state ($\epsilon=0.153$eV from GGA-based DFT calculation) is exactly at the Fermi level, which is consistent with the results of the spectral function. From Fig. S3(d), we can see that at $\omega=0$, the imaginary part of the self-energy for both orbitals is nonzero at $\beta=50$ ($\sim$ 230K), indicating that the DMFT result from GGA functional corresponds to a frozen moment phase Werner et al. (2008) at this temperature, consistent with previous calculations Ouyang et al. (2024). However, this differs from the result obtained from the HSE06 functional in the main text.
@@
The DFT+U is another convenient approach Martin (2020). However, DFT+U only shifts the crystal field splitting without renormalizing the hopping parameters or the bandwidths. In previous works, it was proposed that DFT+U (using GGA functional) can capture the ARPES measurements of the band dispersion in ambient pressure Yang et al. (2024). We also calculate the band structure of the LP phase using the simplified rotation invariant approach to the DFT+U introduced by Dudarev $et$ $al.$ Dudarev et al. (1998) based on GGA functional. As shown in Fig. S3(b), the DFT+U results with $U$=4eV show significant deviation from the ARPES measurements, particularly in terms of the renormalization effect, despite the band positions being nearly correct.
@@
Figure S3: (a) The spectral function $A(k,\omega)$ calculated using DMFT at $U$ = $5.5$ eV, $J_{H}$ = $0.65$ eV, and $\beta=50$ based on TB bands fitted from DFT bands of ordinary GGA functional. The parameters used here are same as the main text. (b) The DFT+U band dispersion using GGA functional at $U=4$ eV. The calculated band dispersion cannot describe the experiment data (red crosses) owing to the band position being incorrect in (a) and missing additional renormalization effects in (b). (c) The real part of the electron self-energy from DMFT calculation. (d) The negative imaginary part of the electron self-energy from DMFT calculation.
VI.3 Hamiltonian and Tight-binding parameters
@@
VI.4 HSE06 band structure of HP phase
We also calculated the band structure and Fermi surface of the HP phase using the HSE06 method, as shown in Fig. S5. As can be seen, due to the doubling of the first Brillouin zone, the top of the $d_{z}$ bonding state, which is located at the $\Gamma$ point in the LP phase, is unfolded back to the M point in the HP phase. However, the most important point is that $d_{z}$ bonding state remains fully occupied.
Figure S5: (a) The HSE06 band structure of HP phase. (b) The HSE06 Fermi surface of HP phase.